In a normal distribution, 69% of the distribution is less than 28 and 90% is less than 35. Find the mean


and standard deviation of the distribution

1
Expert's answer
2022-02-02T16:22:07-0500

We are given that,

"p(Z\lt {28-\mu\over \sigma})=0.69" and "p(Z\lt {35-\mu\over \sigma})=0.90"

So,

"p(Z\lt {28-\mu\over \sigma})=0.69\implies \phi(0.69)={28-\mu\over \sigma}". From the standard normal tables, we have that,

"\phi(0.69)=0.4958503\implies {28-\mu\over \sigma}=0.4958503....(1)"  


and

"p(Z\lt {35-\mu\over \sigma})=0.90\implies \phi(0.90)={35-\mu\over \sigma}". From the standard normal tables, we have that "\phi(0.90)=1.281552" . This implies that,   

"{35-\mu\over \sigma}=1.281552.......(2)"

We shall find the values of "\mu" and "\sigma" by solving equations 1 and 2.

From equation 1,

"{28-\mu\over \sigma}=0.4958503\implies 28-\mu=0.4958503\sigma\implies \sigma={28-\mu\over 0.4958503}.....(3)"

Substituting for the value of "\sigma" from equation 3 in equation 2 we get,

"0.4958503(35-\mu)=1.281552(28-\mu)"

Simplifying this we have,

"17.3547605-0.4958503\mu=35.883456-1.281552\mu\implies0.7857017\mu=18.5286955\implies \mu=23.5823538" From equation 3, we can find the value of "\sigma" as,

"\sigma={28-23.5823538\over 0.4958503}=8.90923369"

Therefore, the values for "\mu" and "\sigma" are 23.58 and 8.91 respectively. Both rounded off to 2 decimal places.